OpenStudy (anonymous):

find the value of x that makes these fractions equivalent x/5=42/70

OpenStudy (amistre64):

multiply both sides by 5

OpenStudy (anonymous):

3/5=42/70

OpenStudy (hoblos):

x=3

OpenStudy (amistre64):

if my head math is right .... we get 42/14 = 21/7 = 3

OpenStudy (anonymous):

x=3 sure